# Problem Space Transformations for Out-of-Distribution Generalisation in Behavioural Cloning

### v2

@@ Abstract
The combination of behavioural cloning and neural networks has driven significant progress in robotic manipulation. As these algorithms may require a large number of demonstrations for each task of interest, they remain fundamentally inefficient in complex scenarios, in which finite datasets can hardly cover the state space. One of the remaining challenges is thus out-of-distribution (OOD) generalisation, i.e. the ability to predict correct actions for states with a low likelihood with respect to the state occupancy induced by the dataset. This issue is aggravated when the system to control is treated as a black-box, ignoring its physical properties. This work characterises widespread properties of robotic manipulation, specifically pose equivariance and locality. We investigate the effect of the choice of problem space on OOD performance of BC policies and how transformations arising from characteristic properties of manipulation could be employed for its improvement. We empirically demonstrate that these transformations allow behaviour cloning policies, using either standard MLP-based one-step action prediction or diffusion-based action-sequence prediction, to generalise better to OOD problem instances.
@@ I INTRODUCTION
The aim of this work is to investigate the effect of the choice of the problem space on the ability of a BC policy to generalize OOD. The focus is on robotic manipulation tasks with rigid bodies and low-dimensional state information. Therefore, practical assumptions on object-centric manipulation tasks are highlighted and leveraged to explore a family of associated problem space transformations. We observe that these transformations are a crucial design component for learning-based control of manipulators, and enable policies learned through BC to perform well on OOD states.
We present three main contributions:
  • we determine properties underlying practical manipulation problems;
@@
  • we provide experimental results demonstrating that the choice of problem space transformation significantly impacts the ability of OOD generalisation for three robotic manipulation tasks.
@@ II Related Work
Past work has explored exploiting different invariances and equivariances in robot manipulation learning. In [22], learnt $\mathrm{SE}(3)$-equivariant object representations, which have an object point cloud as input, are deployed to enable a pick and place system that requires only a few demonstrations. In [23], $\mathrm{SIM}(3)$ equivariance (which additionally to $\mathrm{SE}(3)$ includes scale equivariance) is embedded in the object representation learning module (again from point clouds) as well as in the BC policy network module. The work in [24] takes a similar approach, though adding a $\mathrm{SIM}(3)$-equivariant Diffusion Policy as the BC policy module. Recent work additionally explores the usefulness of the locality of manipulation problems to increase sample efficiency by predicting actions as displacements to points in the scene point cloud [25]. All of the above works assume that the scene entities are sensed as point clouds. Some past work also looks at the benefits of introducing $\mathrm{SO}(2)$-equivariance to online [26] and offline RL [27] where the scene entity poses are assumed to be available.
@@ IV-B Applying Assumptions in Problem Space
[…]
@@ V Experimental Results
In this section we want to understand how BC policies perform in-distribution and OOD, in the baseline problem space $\mathcal{P}$, and in those defined by the proposed transformations. We consider three tasks which fulfil our assumptions on the problem-type to evaluate the effect of the transformations on OOD generalisation.
Figure 5: A rendering of the three environments considered in our evaluation.
@@ VI-A Implementation Details
We train the MLP models using PyTorch. The policy is implemented as a deterministic MLP with ReLU activation functions. We use dropout and $L_{2}$ regularisaition. The models are trained using mini random batches and all data is standardised to zero mean and one standard deviation before being passed to the neural network. The loss function is the 2-norm between policy prediction and dataset sample. The Adam optimizer is used to update the weights. Training time for all tasks is in the order of minutes, while evaluation in simulation is on the order of 10 minutes. The hyperparameters used in training the models is shown in Table I.
TABLE I: Training hyperparameters used when training MLP BC policy for the different tasks.
@@
We use a PyTorch implementation of Diffusion Policy based on the openly available implementation provided by the authors. We follow the same approach to training the models as described in [3] using the transformer-based variant. Training time for all tasks with our implementation is on the order of 10 hours, though significant speed up would be possible with a more efficient implementation of the transformations. The hyperparameters used in training are shown in Table II. The table differentiates between the number of predicted action steps $T_{A_{p}}$ and the number of action steps which are then executed before policy inference is performed again $T_{A_{e}}$.
TABLE II: Training hyperparameters used when training Diffusion Policy for the different tasks.
Task PickPlace Assembly PushT
$T_{X}$ 2 2 2
$T_{A_{p}}$ 10 10 16
$T_{A_{e}}$ 8 8 8
Layers 8 8 8
Learning Rate 1e-4 1e-4 1e-4
Batch Size 256 256 256
Epochs 5010 1800 5010
Attn. Dropout 0.3 0.3 0.01
@@ VI-B Task and Experiment Details
TABLE III: Summary of task and experiment relevant parameters.
Task PickPlace Assembly PushT
# Demonstrations 100 200 200
Collected by Scripted Policy Scripted Policy Human Demonstrator
# In-dist. evaluation episodes 100 100 100
# OOD evaluation episodes 450 450 400
In-dist. manifold obj. pos. $x_{1}\in[-0.15,0.15]$m, $x_{1}\in[-0.15,0.15]$m, $r\in[32,180]$px,
$x_{2}\in[0.48,0.62]$m $x_{2}\in[0.43,0.67]$m $\theta\in[0,2\pi]$rad
OOD manifold obj. pos. $x_{1}\in[-0.465,0.465]$m, $x_{1}\in[-0.465,0.465]$m, $r\in[180,534]$px,
$x_{2}\in[0.4,0.7]$m $x_{2}\in[0.38,0.72]$m $\theta\in[0,2\pi]$rad
$\lambda$ (for $\mathcal{T}_{2}$) 0.1m 0.2m 150px
# Experiment Seeds MLP 8 8 8
# Experiment Seeds DP 5 5 5
Reward Distance to target See implementation [35] Distance to target
Action Space [EE Pos., Gripper Force] [EE Pos., 1D EE orientation, Gripper Force] [EE Pos.]
The main task and experiment specific parameters are summarised in Table III. The specified in-distribution and OOD manifolds relate to the position of the relevant object of the task. For PickPlace and Assembly the in-distribution manifold is a rectangle and the OOD manifold is a larger rectangle excluding the in-distribution manifold. The parameters of these rectangles are provided in Cartesian coordinates in the table. For PushT, they are sampled from a torus with inner and outer radius in both cases, the parameters in the table are provided in cylindrical coordinates. In all cases the end-effector is initialised at the same pose. For Assembly and PushT the orientation of the relevant object is randomised in $[0,2\pi]$ in training and evaluation, but for PickPlace the initial orientation of the object is constant.

### v3

@@ Abstract
The combination of behavioural cloning and neural networks has driven significant progress in robotic manipulation. As these algorithms may require a large number of demonstrations for each task of interest, they remain fundamentally inefficient in complex scenarios, in which finite datasets can hardly cover the state space. One of the remaining challenges is thus out-of-distribution (OOD) generalisation, i.e. the ability to predict correct actions for states with a low likelihood with respect to the state occupancy induced by the dataset. This issue is aggravated when the system to control is treated as a black-box, ignoring its physical properties. This work highlights widespread properties of robotic manipulation, specifically pose equivariance and locality. We investigate the effect of the choice of problem space on OOD performance of BC policies and how transformations arising from characteristic properties of manipulation can be employed for its improvement. Through controlled, simulated and real-world experiments, we empirically demonstrate that these transformations allow behaviour cloning policies, using either standard MLP-based one-step action prediction or diffusion-based action-sequence prediction, to generalise better to certain OOD problem instances. Code is available at https://github.com/kirandoshi/pst_ood_gen.
@@ I INTRODUCTION
The aim of this work is to investigate the effect of the choice of the problem space on the ability of a BC policy to generalize OOD. The focus is on robotic manipulation tasks with rigid bodies and low-dimensional state information. Therefore, practical assumptions on object-centric manipulation tasks are highlighted and leveraged to explore a family of associated problem space transformations. We observe that these transformations are a crucial design component for learning-based control of manipulators, and enable policies learned through BC to perform well on OOD states. Our focus is not that of solving OOD generalisation as a whole, but rather to carefully evaluate how careful transformation of the problem space in a common setup impacts generalization.
We thus present three main contributions:
  • we formalise the OOD problem in the context of manipulation;
@@
  • we provide experimental results demonstrating that the choice of problem space transformation significantly impacts the ability of OOD generalisation for three simulated robotic manipulation tasks and a real-world pushing task.
@@ II Related Work
Past work has explored exploiting different invariances and equivariances in robot manipulation learning. In [25], learnt $\mathrm{SE}(3)$-equivariant object representations, which have an object point cloud as input, are deployed to enable a pick and place system that requires only a few demonstrations. In [31], $\mathrm{SIM}(3)$ equivariance (which additionally to $\mathrm{SE}(3)$ includes scale equivariance) is embedded in the object representation learning module (again from point clouds) as well as in the BC policy network module. The work in [30] takes a similar approach, though adding a $\mathrm{SIM}(3)$-equivariant Diffusion Policy as the BC policy module. Recent work additionally explores the usefulness of the locality of manipulation problems to increase sample efficiency by predicting actions as displacements to points in the scene point cloud [34] (thus assuming 3D sensing) or by leveraging gaze data [26]. Some past work also looks at the benefits of introducing $\mathrm{SO}(2)$-equivariance to online [28] and offline RL [27] where the scene entity poses are assumed to be available.
@@ IV-B Applying Assumptions in Problem Space
The first transformation we consider encodes $\mathrm{SE}(n)$ equivariance to changes of the entities’ poses w.r.t. to a fixed world frame $\mathcal{W}$, where $\mathcal{W}$ measures the state values in a Cartesian coordinate system with fixed origin, denoted as $W$ (e.g., at the base of the robot arm). A state $\mathbf{x}\in\mathcal{X}$ can be expressed as $\mathbf{x}=[\,\prescript{}{\mathcal{W}}{\mathbf{x}_{r}},\prescript{}{\mathcal{W}}{\mathbf{x}_{o,1}},...,\prescript{}{\mathcal{W}}{\mathbf{x}_{o,N_{O}}}\,]$, where the prescript denotes the frame in which the state is measured. Actions are expressed analogously $\mathbf{a}=[\prescript{}{\mathcal{W}}{\mathbf{a}_{r}}]$. We propose to transform $\mathcal{X}$ to a frame $\mathcal{E}$ matching the position and the orientation of the end-effector. This induces a transformed state $\mathcal{T}_{1}(\mathbf{x})=[\,\prescript{}{\mathcal{E}}{\mathbf{x}_{r}},\prescript{}{\mathcal{E}}{\mathbf{x}_{o,1}},...,\prescript{}{\mathcal{E}}{\mathbf{x}_{o,N_{O}}}\,]$. Any action  [^3] This transformation can be applied no matter if the action is supplied as a next EE position, an offset to the current position or as a velocity, though the exact transformation varies. $\mathbf{a}\in\mathcal{A}$ is also transformed accordingly: $\mathcal{T}_{1}(\mathbf{a})=[\prescript{}{\mathcal{E}}{\mathbf{a}_{r}}]$. This ensures that interactions between end-effector and entities may have the same representation, regardless of poses in the fixed coordinate frame $\mathcal{W}$. In turn, this can increase the density of the occupancy over the transformed state space, effectively enlarging the in-distribution manifold. An important aside to $\mathcal{T}_{1}$: For some MDP modelling choices, an additional entity corresponding to a fixed point (usually the target position) in $\mathcal{W}$ needs to be added to $\mathcal{X}$ when transforming the problem space to frame $\mathcal{E}$ such that the demonstrator’s policy remains recoverable.
@@ V Experimental Results
In this section we want to understand how BC policies perform in-distribution and OOD, in the baseline problem space $\mathcal{P}$, and in those defined by the proposed transformations. We first consider three simulated tasks which fulfil our assumptions on the problem-type to evaluate the effect of the transformations on OOD generalisation. While we focus on simulated experiments as they allow better control over experimental conditions, we complete our empirical evaluation on a real-world task in Section V-E, which confirms simulated results on hardware.
Figure 5: A rendering of the three simulation environments considered in our evaluation.
@@ V-D Ablations
V-E Real-World Evaluation
We finally evaluate our approach on a real-world robotic hardware setup to validate the results obtained in simulation. Experiments are conducted on a real-world variant of the PushT task, which differs from the simulated environment and poses additional challenges due to sensing noise and actuation errors.
The task is executed on a single-arm ALOHA robot platform [35]. Training data is collected via teleoperation by a single expert demonstrator. Consistent with the simulation experiments, the policy receives state-based observations as input. The pose of the T-shaped object is estimated using an Apriltag-based tracking system with two calibrated RGB webcams. An overview of the hardware setup is shown in Figure 7.
For the real-world evaluation, we exclusively consider behaviour cloning (BC) with a diffusion policy, due to higher reported performance on hardware [9]. OOD performance results on the physical robot are summarized in Table I. For each policy, we execute 15 rollouts from uniformly sampled initial states, drawn from both an in-distribution (ID) set and an out-of-distribution (OOD) set. The object is initialised in as similar a location as possible across all compared transformation to ensure fair comparison. The first metric we show is the distance from the target the T object should be pushed towards for all three transformations, a lower number is better. The second metric shows how much closer to the goal the object is at termination compared to its initial position, as despite best efforts the exact initial state locations varies between rollouts. The third metric represents the same number as the second metric, though it shows the relative distance moved towards the goal.
Consistent with the simulation results, we observe a clear performance improvement when using $\mathcal{T}_{2}$ compared to the baseline problem space, demonstrating that the benefits of the problem space transformation apply to real-world robotic execution too. Though, in contrast with simulation results, we see that the performance of $\mathcal{T}_{1}$ is reduced compared to the baseline, which in turn means that the increase in performance of $\mathcal{T}_{2}$ can be traced back to the assumption of locality.
TABLE I: Real-World Evaluation
Policy Baseline $\mathcal{T}_{1}$ $\mathcal{T}_{2}$
Mean dist. at termination [m] 0.1977 0.2075 0.1853
Mean dist. moved towards goal [m] 0.0387 0.0292 0.0501
Mean rel. dist. moved towards goal [-] 16.38% 12.32% 21.29%
Figure 7: The real-world one-arm setup (using ALOHA [35]) used to collect teleoperation data and to rollout trained policy (wrist camera is unused).
@@ VI-A Implementation Details
We train the MLP models using PyTorch. The policy is implemented as a deterministic MLP with ReLU activation functions. We use dropout and $L_{2}$ regularisaition. The models are trained using mini random batches and all data is standardised to zero mean and one standard deviation before being passed to the neural network. The loss function is the 2-norm between policy prediction and dataset sample. The Adam optimizer is used to update the weights. Training time for all tasks is in the order of minutes, while evaluation in simulation is on the order of 10 minutes. The hyperparameters used in training the models is shown in Table II.
TABLE II: Training hyperparameters used when training MLP BC policy for the different tasks.
@@
We use a PyTorch implementation of Diffusion Policy based on the openly available implementation provided by the authors. We follow the same approach to training the models as described in [9] using the transformer-based variant. Training time for all tasks with our implementation is on the order of 10 hours, though significant speed up would be possible with a more efficient implementation of the transformations. The hyperparameters used in training are shown in Table III. The table differentiates between the number of predicted action steps $T_{A_{p}}$ and the number of action steps which are then executed before policy inference is performed again $T_{A_{e}}$. To deploy the diffusion policy on a real system we follow [9] by using DDIM noise scheduling and reducing the number of diffusion steps at inference to 16 (100 steps in training) compared with using DDPM noise scheduling in training and testing at 100 diffusion steps for the simulation tasks.
TABLE III: Training hyperparameters used when training Diffusion Policy for the different tasks.
Task PickPlace Assembly PushT PushT
Real
$T_{X}$ 2 2 2 2
$T_{A_{p}}$ 10 10 16 16
$T_{A_{e}}$ 8 8 8 6
Layers 8 8 8 8
Learning Rate 1e-4 1e-4 1e-4 1e-4
Batch Size 256 256 256 256
Epochs 5010 1800 5010 1750
Attn. Dropout 0.3 0.3 0.01 0.1
@@ VI-B Task and Experiment Details
TABLE IV: Summary of task and experiment relevant parameters.
Task PickPlace Assembly PushT PushT Real
# Demonstrations 100 200 200 200
Collected by Scripted Policy Scripted Policy Human Demonstrator Human Demonstrator
# In-dist. evaluation episodes 100 100 100 15
# OOD evaluation episodes 450 450 400 15
In-dist. manifold obj. pos. $x_{1}\in[-0.15,0.15]$m, $x_{1}\in[-0.15,0.15]$m, $r\in[32,180]$px, $x_{1}\in[0.31,0.61]$m
$x_{2}\in[0.48,0.62]$m $x_{2}\in[0.43,0.67]$m $\theta\in[0,2\pi]$rad $x_{2}\in[-0.15,0.15]$m
OOD manifold obj. pos. $x_{1}\in[-0.465,0.465]$m, $x_{1}\in[-0.465,0.465]$m, $r\in[180,534]$px, $x_{1}\in[0.2,0.71]$m
$x_{2}\in[0.4,0.7]$m $x_{2}\in[0.38,0.72]$m $\theta\in[0,2\pi]$rad $x_{2}\in[-0.225,0.225]$m
$\lambda$ (for $\mathcal{T}_{2}$) 0.1m 0.2m 150px 0.225m
# Experiment Seeds MLP 8 8 8 -
# Experiment Seeds DP 5 5 5 1
Reward Distance to target See implementation [32] Distance to target Distance to target
Action Space [EE Pos., Gripper Force] [EE Pos., 1D EE Orientation, Gripper Force] [EE Pos. (2D)] [EE Pos., EE Orient. (quater- nion), Norm. Gripper Pos.]
The main task and experiment specific parameters are summarised in Table IV. The specified in-distribution and OOD manifolds relate to the position of the relevant object of the task. For PickPlace, Assembly and PushT Real the in-distribution manifold is a rectangle and the OOD manifold is a larger rectangle excluding the in-distribution manifold. The parameters of these rectangles are provided in Cartesian coordinates in the table. For PushT, they are sampled from a torus with inner and outer radius in both cases, the parameters in the table are provided in cylindrical coordinates. In all cases the end-effector is initialised at the same pose. For Assembly, PushT and PushT Real the orientation of the relevant object is randomised in $0,2\pi]$ in training and evaluation, but for PickPlace the initial orientation of the object is constant.